OpenStudy (anonymous):

x + 4 2 ----- ÷ --- x^2 x

OpenStudy (espex):

Recall that dividing fractions can be rewritten as multiplying the inverse, so \[\frac{x+4}{x^2} \div \frac{2}{x}\] can be written as \[\frac{x+4}{x^2}\dot{}\frac{x}{2}\] where you only need to multiply numerators, denominators, and simplify if you can.